OpenStudy (republican31):

Solving Rational Equations x+4/5x+6 = 6/x+3 Solutions are x={1,-24} I think is what the lesson said, but I can't get to that point. I don't understand.

ganeshie8 (ganeshie8):

Have you tried cross multiplying ?

ganeshie8 (ganeshie8):

\[\large \dfrac{x+4}{5x+6} = \dfrac{6}{x+3}\] implies \[\large (x+4)(x+3) = (6)(5x+6)\]

ganeshie8 (ganeshie8):

distribute and see what you can do

OpenStudy (republican31):

\[x^2+7x+12=30x+36\] Now what?

ganeshie8 (ganeshie8):

Perfect! collect like terms and make it look like a nice quadratic

ganeshie8 (ganeshie8):

subtract 30x both sides then subtract 36 both sides so that the right hand side becomes 0

OpenStudy (republican31):

\[x^2-23x-24\] Ok I see now thanks!

ganeshie8 (ganeshie8):

Np :) rest should be easy if you're familiar with solving quadratics

OpenStudy (republican31):

Thank you so much I got it now!!!

ganeshie8 (ganeshie8):

glad to hear, yw !